# Targeted Immunisation Thresholds for the Contact Process on Power-Law Trees

### v1

@@ 6.2 Small truncation threshold
where the final line follows from recalling the $\gamma$ asymptotic in (15), that $\lambda|\partial S|\geq\frac{1}{8\lambda^{2}}$ and that $\frac{1}{8}\log\left(7\cdot(3/2)^{-5}\right)<-\frac{1}{100}$. ∎
Lemma 6.5.
Assuming $\lambda$ is sufficiently small, on the tree $\mathscr{T}(S,B)$ we have
$\mathbb{E}_{\mathbbm{1}_{S}}\left(\int_{0}^{\infty}\mathbbm{1}_{\xi_{t}(S)>0}\,{\rm d}t\right)\leq 3e^{200\lambda^{2}|\partial S|}.$
Proof.
As we start the process at the infection $\mathbbm{1}_{S}$, we have $H_{0}=0$. Suppose for now that we keep $H_{t}<20\lambda|\partial S|$ for $t$ long enough (Lemma 6.4 will later give us this control). We also have an infection process on $S$
$U_{t}=\sum_{v\in S}\xi_{t}(v)$
which, while the external infection process has $H_{t}\leq 20\lambda|\partial S|$, has
$\frac{\mathbb{E}\left({\rm d}U_{t}\right)}{{\rm d}t}\leq\lambda H_{t}-U_{t}+\lambda|S|U_{t}\leq 20\lambda^{2}|\partial S|-\frac{1}{2}U_{t},$ (16)
using that for small $\lambda$ we have $2\lambda|S|\leq\frac{2\delta}{1000\rho}\lambda\lceil\delta\log(1/\lambda)\rceil\leq 1$. Further though, this upper bound is just a birth-death chain on $\{0\}\cup[|S|]$ which by detailed balance has a stationary distribution $\pi$ with
$\forall k\in[|S|]\quad 40\lambda^{2}|\partial S|\pi(k-1)=k\pi(k)\implies\frac{\pi(k)}{\pi(0)}=\frac{\left(40\lambda^{2}|\partial S|\right)^{k}}{k!}$
$\pi(0)\frac{\left(40\lambda^{2}|\partial S|\right)^{k}}{2(k-1)!}$
$\frac{1}{\pi(0)}\sum_{k=1}^{|S|}\frac{2(k-1)!}{\left(40\lambda^{2}|\partial S|\right)^{k}}$ $=\sum_{k=1}^{|S|}\frac{2(k-1)!}{\left(40\lambda^{2}|\partial S|\right)^{k}}\sum_{j=1}^{|S|}\frac{\left(40\lambda^{2}|\partial S|\right)^{j}}{j!}$
$\leq|S|\frac{2(|S|-1)!}{\left(40\lambda^{2}|\partial S|\right)^{|S|}}\sum_{j=1}^{\infty}\frac{\left(40\lambda^{2}|\partial S|\right)^{j}}{j!}\leq\frac{2(|S|)!e^{40\lambda^{2}|\partial S|}}{\left(40\lambda^{2}|\partial S|\right)^{|S|}}.$
Then, use that $|\partial S|\geq\frac{|S|}{10\lambda^{2}}$ and $n!\leq\frac{n^{n+1}}{e^{n-1}}$ to bound this hitting time by
$\frac{2(|S|)!e^{40\lambda^{2}|\partial S|}}{\left(40\lambda^{2}|\partial S|\right)^{|S|}}\leq\frac{2(|S|)!e^{40\lambda^{2}|\partial S|}}{\left(4|S|\right)^{|S|}}\leq 2|S|e{(4e)}^{-|S|}e^{40\lambda^{2}|\partial S|}\leq e^{40\lambda^{2}|\partial S|}.$
Once $S$ recovers at some time
$\tau_{1}:=\inf_{t>0}\left\{U_{t}=0\right\},$
we check if $H$ goes to 00 without ever reinfecting the tree $S$ when, as previously mentioned, it recovers while $H_{\tau_{1}}<20\lambda|\partial S|$. Conditionally on $\xi_{\tau_{1}}$, define $\tau_{S}$ the first time after $\tau_{1}$ when some vertex of $S$ is reinfected, and observe that $(H_{t\wedge\tau_{S}}e^{(t\wedge\tau_{S})/4})_{t\geq\tau_{1}}$ is a supermartingale and so
$\mathbb{P}\left(\int_{\tau_{1}}^{\tau_{S}}H_{t}>160\lambda|\partial S|\,\Big{|}\,\xi_{\tau_{1}}\right)\leq\frac{\mathbb{E}\left(\int_{\tau_{1}}^{\tau_{S}}H_{t}{\rm d}t\,\Big{|}\,\xi_{\tau_{1}}\right)}{160\lambda|\partial S|}\leq\frac{20\lambda|\partial S|\int_{0}^{\infty}e^{-t/4}{\rm d}t}{160\lambda|\partial S|}=\frac{1}{2}.$
If this limit is not exceeded, the total rate of reinfection of $S$ is then bounded by $\lambda\cdot 160\lambda|\partial S|$, and so we have found that a recovery of $S$ is permanent with some probability $p$ where $p\geq\frac{1}{2}e^{-160\lambda^{2}|\partial S|}$. For ease of calculation (we would like exact control on the probability of a true recovery), additionally flip a biased coin at time $\tau_{1}$ which is heads with probability $q$ satisfying
$pq=\frac{1}{2}e^{-160\lambda^{2}|\partial S|}.$
If the coin flip lands tails, immediately fully infect $S$ at the time $\tau_{S}=\tau_{1}$ – this guarantees $H_{\tau_{S}}\leq 20\lambda|\partial S|$. Otherwise, the supermartingale property also yields
$\operatorname{\mathbb{E}}[H_{\tau_{S}}\mathbbm{1}_{\tau_{S}<\infty}\,\Big{|}\,\xi_{\tau_{1}}]\leq 20\lambda|\partial S|,$
and after conditioning on the highly probable reinfection of $S$, we still have
$\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}[H_{\tau_{S}}\,\big{|}\,\tau_{S}<\infty]\,\Big{|}\,\xi_{\tau_{1}}\right]\leq\frac{20\lambda|\partial S|}{1-\frac{1}{2}e^{-160\lambda^{2}|\partial S|}}\leq 21\lambda|\partial S|$
which is always true when $\lambda^{2}|\partial S|\geq 1/8$, as is necessary for $S$ to contain any big vertices.
In Lemma 5.1 we use a supermartingale argument to control on the hitting time of $[0,8\lambda|\partial S|]$ from a random initial point using Jensen’s inequality. Rather than repeat these arguments, we will just take the conclusion of (10) to bound the expected hitting time of $[0,8\lambda|\partial S|]$ by
$5348\log\frac{4}{3}+16\log\left(1+\frac{21\lambda|\partial S|}{2\lambda|\partial S|}\right)<$ 53 . (17)
If $(H_{t})$ instead exceeds $20\lambda|\partial S|$ while $S$ is infected, it can be at most $1+20\lambda|\partial S|<21\lambda|\partial S|$ after doing so. Hence from this initial condition we also expect it to hit $[0,8\lambda|\partial S|]$ within time $53$. We can now assemble the upper bound.
  1. (i)
    First we must achieve $H<15\lambda|\partial S|$, which we have seen (in the three possible cases) is never expected to take longer than time $53$.
  2. (ii)
    Wait for hitting of $U=0$ or $H>20\lambda|\partial S|$:
    • By Lemma 6.4 and Markov’s inequality, we will have hitting of $U=0$ before time $7^{\lambda|\partial S|}$, and hitting of $H>20\lambda|\partial S|$ after time $7^{\lambda|\partial S|}$, with probability at least
      $1-\exp\left(-\frac{1}{100\lambda}\right)-\exp\left(\left(40\lambda-\log 7\right)\lambda|\partial S|\right)\geq 1-\exp\left(-\frac{1}{5\lambda}\right);$
    • Otherwise, if we hit $H>20\lambda|\partial S|$, we expect another period of length $53$ to return to $H<15\lambda|\partial S|$, after which we must restart waiting for the recovery of $S$.
    Thus the expected total time in steps (i) and (ii) which we expect to take before finding $U=0$ is bounded by
    $\frac{53+e^{40\lambda^{2}|\partial S|}}{1-e^{-\frac{1}{5\lambda}}}.$
  3. (iii)
    Then a period of healthy $S$, i.e. zero integrand $\mathbbm{1}_{\xi(S)>0}=0$, which will be permanent with probability exactly $\frac{1}{2}e^{-160\lambda^{2}|\partial S|}$. Otherwise, we restart at (i).
Clearly this cycle is made to repeat exactly a geometric number of times and thus
$\mathbb{E}\left(\int_{0}^{\infty}\mathbbm{1}_{\xi_{t}(S)>0}{\rm d}t\Bigg{|}\xi_{0}=\mathbbm{1}_{S}\right)\leq\frac{53+e^{40\lambda^{2}|\partial S|}}{1-e^{-\frac{1}{5\lambda}}}\left(2e^{160\lambda^{2}|\partial S|}\right)$
which gives the claimed result for small $\lambda$. ∎
Corollary 6.6.
On the tree $\mathscr{T}(S,B)$ with the good environment $S$ initially fully infected, fix a target sink $v$ at distance $\lceil\delta\log(1/\lambda)\rceil$ from $S$. The expected number of infections absorbed at this target has
$\operatorname{\mathbb{E}}_{\mathbbm{1}_{S}}\left[\xi^{[v]}_{\infty}(v)\right]\leq e^{-\frac{1}{4}\delta\log^{2}\frac{1}{\lambda}}.$
Proof.
The proof of Corollary 5.2 doesn’t work here as if we put any infection sink adjacent to or even within $O(1)$ distance of a vertex of degree $\mathbbm{t}$, in this regime, we will see local divergence at that star. Hence we must define a faraway sink $w$
$\operatorname{d}(w,S)=\left\lfloor\frac{\delta}{2}\log\frac{1}{\lambda}\right\rfloor$
where $w$ is the unique such vertex disconnecting $v$ and $S$. This distance is attained by some $x\in S$ with $\operatorname{d}(w,x)=\operatorname{d}(w,S)$ and we will bound the infection using the CP-BRW with sinks $\mathcal{V}=\{x,w,v\},$ where the initial infection on $S$ starts to branch and create new types of infections only when hitting $w$ (and not $x$). One may imagine that the initial infection entirely “emanates from $x$” in the branching structure of the CP-BRW types. With this particular initial condition, Inequality (3) should be replaced in this context by
$\operatorname{\mathbb{E}}_{\mathbbm{1}_{S}}[\xi^{[v]}_{\infty}(v)]\leq\operatorname{\mathbb{E}}_{\mathbbm{1}_{S}}[\xi^{[w]}_{\infty}(w)]\sum_{\begin{subarray}{c}k\geq 0,x_{2k+2}=v\\ x_{1}=\ldots=x_{2k+1}=w\\ x_{2}=\ldots=x_{2k}=x\end{subarray}}\prod_{i=2}^{2k+2}m^{\mathcal{V}}(x_{i-1},x_{i}).$ (18)
We upper bound this initial infection by assuming all of $S^{\prime}$ is fully infected, where $S^{\prime}\supset S$ is the minimal extension such that $\operatorname{d}(w,S^{\prime})=1$. That is,
$S^{\prime}:=S\cup[x,w]\setminus\{w\}.$
By Lemma 6.5 we have
$\operatorname{\mathbb{E}}_{\mathbbm{1}_{S}}\left[\xi^{[w]}_{\infty}(w)\right]$ $\leq\lambda\cdot 3e^{200\lambda^{2}|\partial S^{\prime}|}\leq 3\lambda\exp\left({200\lambda^{2}\left(|\partial S|+\frac{1}{8\lambda^{2}}\cdot\frac{\delta}{2}\log\frac{1}{\lambda}\right)}\right)$
$\leq 3\lambda^{1-25\delta/2}e^{200\lambda^{2}|\partial S|}.$
Of course, we have the same bound for $m^{\mathcal{V}}(x,w)$, simply by bounding the infection starting from $x$ infected by that starting with $S^{\prime}$ fully infected. Besides, the usual hostile environment bound [mountford2013metastable, Lemma 5.1(i)] provides the following upper bound available for both $m^{\mathcal{V}}(w,x)$ and $m^{\mathcal{V}}(w,v)$, the mean number of infections sent from $w$ to either $x$ or $v$:
$m^{\mathcal{V}}(w,x)$ $\leq(2\lambda)^{\left\lfloor\frac{\delta}{2}\log\frac{1}{\lambda}\right\rfloor}\leq(2\lambda)^{-1+\frac{\delta}{2}\log\frac{1}{\lambda}}$
$=\exp\left(-\left(\frac{\delta}{2}\log^{2}\frac{1}{\lambda}\right)\left(1-\frac{\delta}{2\log\frac{1}{\lambda}}\right)\left(1-\frac{\log 2}{\log\frac{1}{\lambda}}\right)\right)\leq e^{-\sigma\log^{2}\frac{1}{\lambda}},$
with $\sigma\in(0,\delta/2)$ an arbitrary constant. We now bound the total infection received at $v$ using (18):
$m^{[v]}_{\infty}\leq\sum_{k=1}^{\infty}\left(3\lambda^{1-25\delta/2}e^{200\lambda^{2}|\partial S|}\cdot e^{-\sigma\log^{2}\frac{1}{\lambda}}\right)^{k}.$
Now recalling from (13) the bound $|\partial S|\leq\frac{\delta\log^{2}\frac{1}{\lambda}}{900\lambda^{2}}$ (because we assumed for this corollary that the environment is good), we have:
$e^{200\lambda^{2}|\partial S|-\sigma\log^{2}\frac{1}{\lambda}}\leq e^{\left(\frac{2}{9}\delta-\sigma\right)\log^{2}\frac{1}{\lambda}}.$
Recall that we can take $\frac{2}{9}\delta-\sigma$ arbitrarily close to $\frac{-5}{18}\delta$ and thus bound as claimed
$m^{[v]}_{\infty}\leq\sum_{k=1}^{\infty}\left(3\lambda^{1-25\delta/2}e^{\left(\frac{2}{9}\delta-\sigma\right)\log^{2}\frac{1}{\lambda}}\right)^{k}\leq\sum_{k=1}^{\infty}\left(e^{-\frac{5}{19}\delta\log^{2}\frac{1}{\lambda}}\right)^{k}\leq e^{-\frac{1}{4}\delta\log^{2}\frac{1}{\lambda}}.$
By Corollary 6.6, on a good environment the probability that the infection ever reaches any given vertex $v$ of $S^{+}$ at maximal distance $\lceil\delta\log(1/\lambda)\rceil$ from $S$ is bounded by $e^{-\frac{1}{4}\delta\log^{2}\frac{1}{\lambda}}$. Recalling that this good environment has $|S^{+}|\leq\left\lfloor\frac{\delta}{1000\rho}\right\rfloor\mathbbm{t}^{1+\delta\log\mu}$, by the union bound the survival probability of the infection on this good environment is bounded by
$\left\lfloor\frac{\delta}{1000\rho}\right\rfloor\mathbbm{t}^{1+\delta\log\mu}e^{-\frac{1}{4}\delta\log^{2}\frac{1}{\lambda}}$
@@ Appendix A A lemma on the contact process on star graphs
Consider the contact process $(\xi_{t})$ with infection rate $\lambda<1$ on the star graph $S_{k}$. Assume that $\rho$ is infected at time 0, and write $T=\int_{0}^{\infty}\mathbbm{1}_{\rho\in\xi_{t}}\operatorname{d}t$ the total time the center is infected before the extinction of the epidemic on $S_{k}$. Then the probability of $\{T\leq e^{\varepsilon\lambda^{2}k}\}$ is bounded below by $c/\lambda^{2}k$, for $\varepsilon>0$ and $c>0$ universal constants (i.e independent of $\lambda<1$ and $k$).
@@
This result states that the probability that the contact process fails to survive for an exponentially long time (at the center of the star) is bounded above by $c/\lambda^{2}k$. This is to be compared with [nguyen2022subcritical, Lemma 3.4], with mainly two differences:
@@
    To prove this, just argue that conditionally on the first recovery time $R$ of the root, each leaf has independently probability $(1-e^{-\lambda R})\left(\frac{\lambda}{1+\lambda}\right)\leq\lambda^{2}R$ to be infected by the root before time $R$ and then send an infection back before its own recovery. Thus, conditionally on $R\leq 1/2\lambda^{2}k$, the probability that the center is reinfected by neighbour is at most $1/2$. We conclude as $R$ is a standard exponential random variable.
@@
$\mathbb{P}(N_{R}<\lfloor\lambda kx\rfloor)\leq ax\qquad\forall x\leq 1/a.$
@@
We have to prove that for any positive integer $i$ smaller than $\lambda kx_{0}$, we have $\mathbb{P}(N_{R}<i)\leq\frac{ai}{1+\lambda k}$. Call $M$ the total number of leaves that receive an infection on the time interval $[0,R]$, but are not necessarily infected anymore at time $R$, so $M\geq N_{R}$. We first treat separately the cases $R\geq 1$ and $\{R<1,M\geq k/2\}$, on which $N_{R}$ is very unlikely to be small. On the event $\{R\geq 1\},$ we simply look at infections of the root on the time interval $I=[R-1,R]$ and argue that each leaf will independently receive an infection on $I$ and experience no recovery on $I$ with probability $(1-e^{-\lambda})e^{-1}\geq\lambda/3$. We conclude with a standard Chernoff bound for Bernoulli random variables
$\mathbb{P}(R\geq 1,N_{R}<i)\leq\mathbb{P}(\operatorname{Bin}(k,\lambda/e)\leq\frac{\lambda k}{4})\leq\exp(-\frac{\lambda k}{32})\leq\frac{a}{1+\lambda k}\leq\frac{ai}{1+\lambda k}$
for $41\leq i\leq\lambda k/$ 4 and $a$ large. On the event $\{R<1,M\geq k/2\}$, each of the $M$ infected leaves has probability $e^{-1}$ to experience no recovery on the whole time interval $[0,1]$ and thus be infected at time $R$. Thus we can lower bound $N_{R}$ by a $\operatorname{Bin}(k/2,e^{-1})$ and conclude similarly. Finally, while the center has not recovered and at least $k/2$ have still not been infected, we infect a new leaf that then stays infected during one unit of time with rate at least $(\lambda k/2)e^{-1}\geq\lambda k/6$. Hence, the probability of the center recovering before $i$ such infection events is bounded by
$i.\frac{1}{1+\lambda k/6}\leq\frac{6i}{\lambda k}.$
@@
[nguyen2022subcritical, Lemmas 3.2 and 3.3] Starting from a configuration with $M$ infected neighbours at time $t$, the central vertex will be infected for at least half of the time-interval $[t,t+1]$ and there will be at least $\frac{\lambda k}{100}$ infected neighbours at time $t+1$ with probability at least
$1-3\exp(-\lambda M/128).$
@@
$\mathbb{P}\left(N_{R+1}\leq\frac{\lambda k}{100}\right)\leq\operatorname{\mathbb{E}}\left[\mathbb{P}\left(N_{R+1}\leq\frac{\lambda k}{100}\Big{|}N_{R}\right)\right]\leq\operatorname{\mathbb{E}}\left[3\exp(-\lambda N_{R}/128)\right],$
where by Lemma A.3 we have that $(1+N_{R})/\lambda k$ is stochastically bounded below by $U$ a uniform random variable on $[0,1/a]$, and thus

### v2

@@ 6.2 Small truncation threshold
where the final line follows from recalling the $\gamma$ asymptotic in (16), that $\lambda|\partial\mathcal{C}|\geq\frac{1}{8\lambda^{2}}$ and that $\frac{1}{8}\log\left(7\cdot(3/2)^{-5}\right)<-\frac{1}{100}$. ∎
Acknowledgements. JF was supported by the Unité de mathématiques pures et appliqués of ENS Lyon, and then by NRDI grant KKP 137490.
[…]
@@ Appendix A A lemma on the contact process on star graphs
Consider the contact process $(\xi_{t})$ with infection rate $\lambda<1$ on the star graph $S_{k}$. Assume that $\rho$ is infected at time 0, and write $T=\int_{0}^{\infty}\mathbbm{1}_{\rho\in\xi_{t}}\operatorname{d}t$ the total time the center is infected before the extinction of the epidemic on $S_{k}$. Then
$\operatorname{\mathbb{P}}\left(T\leq e^{\varepsilon\lambda^{2}k}\right)\leq\frac{c}{\lambda^{2}k},$
for $\varepsilon>0$ and $c>0$ universal constants (i.e independent of $\lambda<1$ and $k$).
@@
This result states that the probability that the contact process fails to survive for an exponentially long time (at the center of the star) is bounded below by $c/\lambda^{2}k$. This is to be compared with [nguyen2022subcritical, Lemma 3.4], with mainly two differences:
@@
    To prove this, just argue that conditionally on the first recovery time $R$ of the root, each leaf has independently probability $(1-e^{-\lambda R})\left(\frac{\lambda}{1+\lambda}\right)\leq\lambda^{2}R$ to be infected by the root before time $R$ and then send an infection back before its own recovery. Thus, conditionally on $R\leq{1}/{2\lambda^{2}k}$, the probability that the center is reinfected by neighbour is at most $\nicefrac{{1}}{{2}}$. We conclude as $R$ is a standard exponential random variable.
@@
$\mathbb{P}(N_{R}<\lfloor\lambda kx\rfloor)\leq ax\qquad\forall x\leq\nicefrac{{1}}{{a}}.$
@@
We have to prove that for any positive integer $i$ smaller than $\lambda kx_{0}$, we have
$\mathbb{P}(N_{R}<i)\leq\frac{ai}{1+\lambda k}.$
Call $M$ the total number of leaves that receive an infection on the time interval $[0,R]$, but are not necessarily infected anymore at time $R$, so $M\geq N_{R}$. We first treat separately the cases $R\geq 1$ and $\{R<1,M\geq\nicefrac{{k}}{{2}}\}$, on which $N_{R}$ is very unlikely to be small.
On the event $\{R\geq 1\},$ we simply look at infections of the root on the time interval $I=[R-1,R]$ and argue that each leaf will independently receive an infection on $I$ and experience no recovery on $I$ with probability $(1-e^{-\lambda})e^{-1}\geq\nicefrac{{\lambda}}{{3}}$. We conclude with a standard Chernoff bound for Bernoulli random variables
$\mathbb{P}(R\geq 1,N_{R}<i)\leq\mathbb{P}\left(\operatorname{Bin}\left(k,\nicefrac{{\lambda}}{{3}}\right)\leq\frac{\lambda k}{4}\right)\leq\exp\left(-\frac{\lambda k}{32}\right)\leq\frac{a}{1+\lambda k}\leq\frac{ai}{1+\lambda k}$
for $41\leq i\leq\lambda k/$ 4 and $a$ large.
On the event $\{R<1,M\geq\nicefrac{{k}}{{2}}\}$, each of the $M$ infected leaves has probability $e^{-1}$ to experience no recovery on the whole time interval $[0,1]$ and thus be infected at time $R$. Thus we can lower bound $N_{R}$ by a $\operatorname{Bin}(\nicefrac{{k}}{{2}},e^{-1})$ and conclude similarly.
Finally, while the center has not recovered and at least $\nicefrac{{k}}{{2}}$ have still not been infected, we infect a new leaf that then stays infected during one unit of time with rate at least $(\lambda k/2)e^{-1}\geq\lambda k/6$. Hence, the probability of the center recovering before $i$ such infection events is bounded by
$i\cdot\frac{1}{1+\lambda k/6}\leq\frac{6i}{\lambda k}.$
@@
[nguyen2022subcritical, Lemmas 3.2 and 3.3] Starting from a configuration with $M$ infected neighbours at time $t$, the central vertex will be infected for at least half of the time-interval $[t,t+1]$ and there will be at least $\lambda k/100$ infected neighbours at time $t+1$ with probability at least
$1-3\exp\left(-\frac{\lambda M}{128}\right).$
@@
$\mathbb{P}\left(N_{R+1}\leq\frac{\lambda k}{100}\right)\leq\operatorname{\mathbb{E}}\left[\mathbb{P}\left(N_{R+1}\leq\frac{\lambda k}{100}\bigg{|}N_{R}\right)\right]\leq\operatorname{\mathbb{E}}\left[3\exp\left(-\frac{\lambda N_{R}}{128}\right)\right],$
where by Lemma A.3 we have that $(1+N_{R})/\lambda k$ is stochastically bounded below by $U$ a uniform random variable on $[0,\nicefrac{{1}}{{a}}]$, and thus